Find three consecutive odd intergers whose sum is 105.
Answer by elima(1433)

Find three consecutive odd intergers whose sum is 105.
The consecutive numbers will be;
x, x+2, x+4
x+(x+2)+(x+4)=105
3x+6=105
3x=105-6
3x=99
x=33
Now that you have the first odd integer, you think you could get the others?
Hope you understand
=)